OpenStudy (anonymous):

f(x) = 5x + 3; g(x) = 6x - 5 Find f/g. 5x + 3/6x - 5 domain {x|x ≠3/5 or 5x + 3/6x - 5 domain {x|x ≠ 5/6 ?

OpenStudy (luigi0210):

Do you know how to find the domain?

OpenStudy (anonymous):

I'm pretty sure it's the second one because the bottom can't equal to 0 right?

OpenStudy (luigi0210):

You are correct sir!

OpenStudy (anonymous):

Thank you! I just don't want to take chances with picking the wrong one XD Could you maybe help me with one more question?

OpenStudy (luigi0210):

Sure.

OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = 8/x^2 + 4 I want to say g(x)= 1/x f(x)= 8/x +4 but it doesn't look right.

OpenStudy (luigi0210):

Okay, there are many possibilities for something like this. I would say the easiest way, since it's a fraction, would be to make the denominator an x. and make g(x) whatever is in the denominator

OpenStudy (luigi0210):

If you plug in what you got we would get this: \[\frac{8}{\frac{1}{x}+4}\]

OpenStudy (luigi0210):

Are you with me so far or did I lose you? :/

OpenStudy (anonymous):

I'm a bit lost XD But does this look alright? I think I got it on my own. g(x)= 1/x^3 f(x)= 8x +4x^3

OpenStudy (luigi0210):

Hm, well plug it in, where ever you see an x for f(x) plug in g(x)

OpenStudy (anonymous):

 Because the x's cancel each other out right?

OpenStudy (luigi0210):

No, you don't keep the x, you plug it INTO the x.

OpenStudy (luigi0210):

If you had this: f(3)=x^2-1 What would you do?

OpenStudy (anonymous):

Oh. Then carry on XD

OpenStudy (anonymous):

3^2-1

OpenStudy (luigi0210):

Exactly, and f(g(x)) is the same concept, expect you're using equations instead of numbers.

OpenStudy (anonymous):

Would it be this simple then? g(x)= x^2 f(x)= 8/x +4

OpenStudy (luigi0210):

 The simplest way to think of it, like I said earlier was to just separate them into two different equations